# Timelike conjugate points in Lorentzian length spaces

### v2

@@ 1 Introduction
Our article is structured as follows: In Section 2, we first recall some basics on Lorentzian pre-length spaces used throughout (Subsection 2.1). We then study geodesics and their convergence properties in Subsection $2.$ Due to the lack of arclength parametrization in the generality that we work with, we need to consider equivalence classes of geodesics up to suitable reparametrizations. To this end, in Appendix A, we give a treatment of the Fréchet distance on such equivalence classes of non-stopping curves on a metric spaces that is suitable for our purposes. Our main result in this context is that the set of (equivalence classes of) causal geodesics $\operatorname{Geo}(X)$ is closed in the set of non-stopping curves with respect to the Fréchet distance (see Proposition 2.26).
@@ 2.1 Preliminaries on Lorentzian pre-length spaces
future-directed timelike (resp. causal) curve (past-directed curves are defined analogously) in a Lorentzian pre-length space $X$ is a nowhere constant map $\gamma\colon I=[a,b]\to X$ (i.e., $\gamma$ is not constant on any non-trivial subinterval of II) that is Lipschitz (w.r.t. $d$) and such that for all $t_{1},t_{2}\in I$ with $t_{1}<t_{2}$ we have $\gamma(t_{1})\ll\gamma(t_{2})$ (resp.) $\gamma(t_{1})\leq\gamma(t_{2})$ (causal curves on unbounded intervals can be defined analogously). Its $\tau$-length is defined as
@@ 2.2 Convergence of geodesics
A future-directed causal curve $\gamma\colon I\to X$ is a geodesic if for each $t\in I$ there exists a neighborhood $[a,b]$ of $t$ (i.e., $a<t<b$, allowing for equality at the endpoints of II) such that $\gamma|_{[a,b]}$ is a maximizer, i.e., $\tau(\gamma(a),\gamma(b))=L_{\tau}(\gamma|_{[a,b]})$.
@@
Let $(X,d,\ll,\leq,\tau)$ be a finitely $\tau$-measurable Lorentzian pre-length space. We denote by $\operatorname{Geo}(X)$ the set of equivalence classes of future-directed causal geodesic segments $\gamma\colon[0,1]\to X$, where two such segments are called equivalent if and only if there exists a strictly increasing and surjective reparametrization $\varphi\colon[0,1]\to[0,1]$ between them. [^6] See Definition $2.$ We denote by $\operatorname{TGeo}(X)\subseteq\operatorname{Geo}(X)$ the subset of equivalence classes of future-directed timelike geodesic segments. For convenience, we write $\gamma\in\operatorname{Geo}(X)$ for both the equivalence class and some representative.
@@
    $\gamma_{n}^{\prime}(0)\to\gamma^{\prime}(0)$ in TMTM.
@@
Clearly, $(ii)$ implies $(i)$. To see that $(iii)$ implies $(ii)$, observe that if $(p_{n},v_{n})\in TM$ with $(p_{n},v_{n})\to(p,v)$, then the fact that $\gamma_{n}(t)\coloneqq\exp_{p_{n}}(tv_{n})$ converges to $\gamma(t)\coloneqq\exp_{p}(tv)$ in $C^{\infty}_{\mathrm{loc}}$ on the maximal domain of $\gamma$ is a consequence of basic ODE theory, see e.g. [12, Cor. 2.6]. We are left with showing that $(i)$ implies $(iii)$. Suppose $\gamma_{n}\colon[a,b]\to M$ converges uniformly to $\gamma\colon[a,b]\to M$. It is well-known that the exponential map $E\colon\mathcal{D}\subseteq TM\to M^{2}$ is a diffeomorphism from a neighborhood of the zero section in TMTM onto a neighborhood of the diagonal in $M^{2}$. We may cover $\gamma$ with finitely many convex normal neighborhoods $U_{i}$ such that $\bigcup_{i}U_{i}^{2}$ is contained in the aforementioned neighborhood of the diagonal. For large enough $n$, the $U_{i}$ cover the corresponding portions of $\gamma_{n}$ due to uniform convergence. But then
@@ 3.2 Families of timelike geodesics
Let $v_{i}=\gamma_{i}^{\prime}(0)\in TM$ and $v=\gamma^{\prime}(0)\in T_{p}M$. By Lemma 2.29, we have $v_{i}\to v$. Consider the map $E:TM\supseteq\mathcal{D}\to M\times M$ defined on a suitable open neighbourhood $\mathcal{D}$ of the 0-section in TMTM by
@@ 3.3 Unreachable and ultimate conjugate points
In [21, Ex. 4.7], an example (due to V. Schroeder) of a geodesic $\sigma\colon[0,1]\to N$ in a complete Riemannian manifold $(N,h)$ (specifically, a two-dimensional ellipsoid) is given, with the following properties: $\sigma$ is the unique geodesic in NN connecting $x=\sigma(0)$ and $y=\sigma(1)$, while $x$ and $y$ are conjugate along $\sigma$. We may “Lorentzify” this example as follows: Let $M\coloneqq\mathbb{R}\times N$ with Lorentzian metric $g=-dt^{2}+h_{x}$. Then $(M,g)$ is globally hyperbolic (each slice $\{t=t_{0}\}$ is a Cauchy surface), and the following correspondences hold (cf., e.g., [18, Ch. 7]): Any geodesic in $(M,g)$ is of the form $s\mapsto(as+b,\beta)$, where $a,b\in\mathbb{R}$ and $\beta$ is a geodesic in $(N,h)$. Jacobi fields in $(M,g)$ take the form $s\mapsto(T(s),J(s))$, where $T$ is affine linear and $J$ is a Jacobi field in $(N,h)$. Thus if we set $\gamma(s)\coloneqq(as,\sigma(s))$ with $\sigma$ as above, then for $a>0$ sufficiently large, $\gamma$ is a timelike geodesic connecting $p=(0,x)$ to $q=(a,y)$. It is the unique timelike geodesic in $(M,g)$ connecting $p$ and $q$, and these points are conjugate along $\gamma$. Since $\sigma$ is minimizing from $x$ to $y$ in NN, $\gamma$ is a timelike maximizer from $p$ to $q$ in $(M,g)$. By Corollary 3.22, $p$ and $q$ are reachable along $\gamma$. Consequently, they provide an example of points along a timelike geodesic that are (by Proposition 3.5) symmetrically (hence ultimate) conjugate, but not unreachable conjugate.
@@ 3.4 Causal cut loci
In the case where there is always a $q^{\prime}$ as in (ii), we take $p$ to be the initial future causal cut point of itself along $\gamma$. We denote the set of initial future causal cut points of $p$ along causal geodesics by $\mathrm{IniCut}_{C}^{+}(p)\subseteq X$, called the initial future causal cut locus of pp. If $X$ is regular, we define the initial future timelike cut locus $\mathrm{IniCut}_{T}^{+}(p)$ and the initial future null cut locus $\mathrm{IniCut}_{N}^{+}(p)$ by replacing all instances of causal with timelike resp. null. It is then clear that $\mathrm{IniCut}_{C}^{+}(p)$ is the disjoint union of $\mathrm{IniCut}_{T}^{+}(p)$ and $\mathrm{IniCut}_{N}^{+}(p)$.
@@
Let $r\geq 0$ be such that all causally related points of distance at most $r$ can be connected by a unique maximizer, let $p\in X$ and $q\in\mathrm{IniCut}_{C}^{+}(p)$. Then by definition, either there exist two maximizers from $p$ to $q$ (in which case $d(p,q)\geq r$), or, arbitrarily close to $q$ there exist cut points of pp. In the latter case, if $d(p,q)<r$, let $\varepsilon>0$ and $\tilde{q}$ a cut point of $p$ so that $d(p,\tilde{q})\leq d(p,q)+\varepsilon<r$, which is a contradiction. So in any case, $d(p,q)\geq r$, and taking the infimum over $q$ and then $p$, as well as the supremum over $r\geq 0$, we get
@@ 4.1 A timelike Rauch comparison theorem
We may assume that $\bar{\gamma}_{n},\bar{\sigma}_{n}$ are parametrized to unit speed. Consider now a breakpoint $p$ of $\bar{\gamma}_{n}$, and for the moment call $\bar{\alpha}_{n}$ and $\bar{\beta}_{n}$ the geodesic sub-segments of $\bar{\gamma}_{n}$ meeting at pp. Then replace $\bar{\beta}_{n}$ by its image under an isometry of $\mathbb{L}^{2}(\kappa)$ whose derivative maps the initial velocity vector of $\bar{\beta}_{n}$ to the terminal velocity vector of $\bar{\alpha}_{n}$, thereby straightening the concatenation of $\bar{\alpha}_{n}$ and $\bar{\beta}_{n}$ into an unbroken geodesic segment. We repeat this procedure at each breakpoint of $\bar{\gamma}_{n}$, and analogously for $\bar{\sigma}_{n}$. Concavity of the stacking implies that these unbroken geodesics will remain within the region bounded by the original stacking and therefore also have to meet. Moreover, since we applied an isometry in each step, the resulting unbroken geodesic segments retain the Lorentzian length of $\bar{\gamma}_{n}$ and $\bar{\sigma}_{n}$, respectively. Consequently, we have found two different geodesic segments in $\mathbb{L}^{2}(\kappa)$ that meet at $\bar{\tau}$-length strictly less than $D_{\kappa}$, a contradiction. ∎
@@ 4.2 Nonexistence of conjugate points under nonpositive timelike curvature
Cover $\gamma([0,1])$ by comparison neighborhoods $V_{1},\dots,V_{N}$ such that for a suitable partition $0=t_{0}<\dots<t_{N}=1$, $\gamma([t_{i},t_{i+1}])\subseteq V_{i+1}$ for all $i=0,\dots,N-1$. By (ii), there exists some $k_{0}$ such that also $\gamma_{p_{k}q_{k}}([t_{i},t_{i+1}])\subseteq V_{i+1}$ for all $k\geq k_{0}$ and all ii. The proof of [7, Prop. 2.3] shows that then each $\gamma_{p_{k}q_{k}}|_{[t_{i},t_{i+1}]}$ as well as $\gamma|_{[t_{i},t_{i+1}]}$ are timelike maximizers. Since $\tau|_{V_{i}\times V_{i}}$ is continuous, we obtain
@@ Appendix A The Fréchet distance between curves
Let $\alpha\colon[0,1]\to[0,1]^{2}$ be a continuous, coordinate-wise non-decreasing map given by $\alpha(t)=(\varphi(t),\psi(t))$ with $\alpha(0)=(0,0)$ and $\alpha(1)=(1,1)$. Then there exists a coordinate-wise non-decreasing and 11-Lipschitz continuous curve $\tilde{\alpha}=(\tilde{\varphi},\tilde{\psi})\colon[0,2]\to[0,1]^{2}$ that has the same image as $\alpha$.
@@
Finally, to show that $\tilde{\alpha}$ is 11-Lipschitz, let $0\leq u<v\leq 2$. Then using monotonicity of $\theta$, $\varphi$ and $\psi$, hence of $\tilde{\varphi}$ and $\tilde{\psi}$,
@@
Let NN be the set of continuous, coordinate-wise non-decreasing maps $\alpha=(\varphi,\psi)\colon[0,1]\to[0,1]^{2}$ with $\alpha(0)=(0,0)$ and $\alpha(1)=(1,1)$. Then $\mathcal{N}\coloneqq\{\alpha([0,1])\mid\alpha\in N\}$ is compact with respect to the Hausdorff distance $d_{H}$.
@@
Denote by $L$ the set of all coordinate-wise non-decreasing 11-Lipschitz maps $\eta\colon[0,2]\to[0,1]^{2}$ with $\eta(0)=(0,0)$ and $\eta(1)=(1,1)$. By the Theorem of Arzelà–Ascoli, $L$ is relatively compact in $C([0,2],[0,1]^{2})$. Since it is clearly also closed with respect to uniform convergence, it is indeed compact in $C([0,2],[0,1]^{2})$.
@@
Since $d_{H}(\iota(\eta),\iota(\zeta))\leq\sup_{u\in[0,2]}\|\eta(u)-\zeta(u)\|_{2}$, it follows that $\iota$ is continuous (indeed, 11-Lipschitz). Thus $\mathcal{L}=\iota(L)$ is compact with respect to $d_{H}$. Finally, by Lemma A.1, $\mathcal{N}=\mathcal{L}$, concluding the proof. ∎
@@ Acknowledgments
This research was funded in part by the Austrian Science Fund (FWF) [Grant DOI 10.55776/EFP6 and 10.55776/J4913]. For open access purposes, the authors have applied a $C$ BY public copyright license to any author accepted manuscript version arising from this submission.

### v3

@@ 1 Introduction
Our article is structured as follows: In Section 2, we first recall some basics on Lorentzian pre-length spaces used throughout (Subsection 2.1). We then study geodesics and their convergence properties in Subsection 2.2. Due to the lack of arclength parametrization in the generality that we work with, we need to consider equivalence classes of geodesics up to suitable reparametrizations. To this end, in Appendix A, we give a treatment of the Fréchet distance on such equivalence classes of non-stopping curves on a metric spaces that is suitable for our purposes. Our main result in this context is that the set of (equivalence classes of) causal geodesics $\operatorname{Geo}(X)$ is closed in the set of non-stopping curves with respect to the Fréchet distance (see Proposition 2.26).
@@ 2.1 Preliminaries on Lorentzian pre-length spaces
future-directed timelike (resp. causal) curve (past-directed curves are defined analogously) in a Lorentzian pre-length space $X$ is a nowhere constant map $\gamma\colon I=[a,b]\to X$ (i.e., $\gamma$ is not constant on any non-trivial subinterval of $I$) that is Lipschitz (w.r.t. $d$) and such that for all $t_{1},t_{2}\in I$ with $t_{1}<t_{2}$ we have $\gamma(t_{1})\ll\gamma(t_{2})$ (resp.) $\gamma(t_{1})\leq\gamma(t_{2})$ (causal curves on unbounded intervals can be defined analogously). Its $\tau$-length is defined as
@@ 2.2 Convergence of geodesics
A future-directed causal curve $\gamma\colon I\to X$ is a geodesic if for each $t\in I$ there exists a neighborhood $[a,b]$ of $t$ (i.e., $a<t<b$, allowing for equality at the endpoints of $I$) such that $\gamma|_{[a,b]}$ is a maximizer, i.e., $\tau(\gamma(a),\gamma(b))=L_{\tau}(\gamma|_{[a,b]})$.
@@
Let $(X,d,\ll,\leq,\tau)$ be a finitely $\tau$-measurable Lorentzian pre-length space. We denote by $\operatorname{Geo}(X)$ the set of equivalence classes of future-directed causal geodesic segments $\gamma\colon[0,1]\to X$, where two such segments are called equivalent if and only if there exists a strictly increasing and surjective reparametrization $\varphi\colon[0,1]\to[0,1]$ between them. [^6] See Definition 2.2. We denote by $\operatorname{TGeo}(X)\subseteq\operatorname{Geo}(X)$ the subset of equivalence classes of future-directed timelike geodesic segments. For convenience, we write $\gamma\in\operatorname{Geo}(X)$ for both the equivalence class and some representative.
@@
    $\gamma_{n}^{\prime}(0)\to\gamma^{\prime}(0)$ in $TM$.
@@
Clearly, $(ii)$ implies $(i)$. To see that $(iii)$ implies $(ii)$, observe that if $(p_{n},v_{n})\in TM$ with $(p_{n},v_{n})\to(p,v)$, then the fact that $\gamma_{n}(t)\coloneqq\exp_{p_{n}}(tv_{n})$ converges to $\gamma(t)\coloneqq\exp_{p}(tv)$ in $C^{\infty}_{\mathrm{loc}}$ on the maximal domain of $\gamma$ is a consequence of basic ODE theory, see e.g. [12, Cor. 2.6]. We are left with showing that $(i)$ implies $(iii)$. Suppose $\gamma_{n}\colon[a,b]\to M$ converges uniformly to $\gamma\colon[a,b]\to M$. It is well-known that the exponential map $E\colon\mathcal{D}\subseteq TM\to M^{2}$ is a diffeomorphism from a neighborhood of the zero section in $TM$ onto a neighborhood of the diagonal in $M^{2}$. We may cover $\gamma$ with finitely many convex normal neighborhoods $U_{i}$ such that $\bigcup_{i}U_{i}^{2}$ is contained in the aforementioned neighborhood of the diagonal. For large enough $n$, the $U_{i}$ cover the corresponding portions of $\gamma_{n}$ due to uniform convergence. But then
@@ 3.2 Families of timelike geodesics
Let $v_{i}=\gamma_{i}^{\prime}(0)\in TM$ and $v=\gamma^{\prime}(0)\in T_{p}M$. By Lemma 2.29, we have $v_{i}\to v$. Consider the map $E:TM\supseteq\mathcal{D}\to M\times M$ defined on a suitable open neighbourhood $\mathcal{D}$ of the 0-section in $TM$ by
@@ 3.3 Unreachable and ultimate conjugate points
In [21, Ex. 4.7], an example (due to V. Schroeder) of a geodesic $\sigma\colon[0,1]\to N$ in a complete Riemannian manifold $(N,h)$ (specifically, a two-dimensional ellipsoid) is given, with the following properties: $\sigma$ is the unique geodesic in $N$ connecting $x=\sigma(0)$ and $y=\sigma(1)$, while $x$ and $y$ are conjugate along $\sigma$. We may “Lorentzify” this example as follows: Let $M\coloneqq\mathbb{R}\times N$ with Lorentzian metric $g=-dt^{2}+h_{x}$. Then $(M,g)$ is globally hyperbolic (each slice $\{t=t_{0}\}$ is a Cauchy surface), and the following correspondences hold (cf., e.g., [18, Ch. 7]): Any geodesic in $(M,g)$ is of the form $s\mapsto(as+b,\beta)$, where $a,b\in\mathbb{R}$ and $\beta$ is a geodesic in $(N,h)$. Jacobi fields in $(M,g)$ take the form $s\mapsto(T(s),J(s))$, where $T$ is affine linear and $J$ is a Jacobi field in $(N,h)$. Thus if we set $\gamma(s)\coloneqq(as,\sigma(s))$ with $\sigma$ as above, then for $a>0$ sufficiently large, $\gamma$ is a timelike geodesic connecting $p=(0,x)$ to $q=(a,y)$. It is the unique timelike geodesic in $(M,g)$ connecting $p$ and $q$, and these points are conjugate along $\gamma$. Since $\sigma$ is minimizing from $x$ to $y$ in $N$, $\gamma$ is a timelike maximizer from $p$ to $q$ in $(M,g)$. By Corollary 3.22, $p$ and $q$ are reachable along $\gamma$. Consequently, they provide an example of points along a timelike geodesic that are (by Proposition 3.5) symmetrically (hence ultimate) conjugate, but not unreachable conjugate.
@@ 3.4 Causal cut loci
In the case where there is always a $q^{\prime}$ as in (ii), we take $p$ to be the initial future causal cut point of itself along $\gamma$. We denote the set of initial future causal cut points of $p$ along causal geodesics by $\mathrm{IniCut}_{C}^{+}(p)\subseteq X$, called the initial future causal cut locus of $p$. If $X$ is regular, we define the initial future timelike cut locus $\mathrm{IniCut}_{T}^{+}(p)$ and the initial future null cut locus $\mathrm{IniCut}_{N}^{+}(p)$ by replacing all instances of causal with timelike resp. null. It is then clear that $\mathrm{IniCut}_{C}^{+}(p)$ is the disjoint union of $\mathrm{IniCut}_{T}^{+}(p)$ and $\mathrm{IniCut}_{N}^{+}(p)$.
@@
Let $r\geq 0$ be such that all causally related points of distance at most $r$ can be connected by a unique maximizer, let $p\in X$ and $q\in\mathrm{IniCut}_{C}^{+}(p)$. Then by definition, either there exist two maximizers from $p$ to $q$ (in which case $d(p,q)\geq r$), or, arbitrarily close to $q$ there exist cut points of $p$. In the latter case, if $d(p,q)<r$, let $\varepsilon>0$ and $\tilde{q}$ a cut point of $p$ so that $d(p,\tilde{q})\leq d(p,q)+\varepsilon<r$, which is a contradiction. So in any case, $d(p,q)\geq r$, and taking the infimum over $q$ and then $p$, as well as the supremum over $r\geq 0$, we get
@@ 4.1 A timelike Rauch comparison theorem
We may assume that $\bar{\gamma}_{n},\bar{\sigma}_{n}$ are parametrized to unit speed. Consider now a breakpoint $p$ of $\bar{\gamma}_{n}$, and for the moment call $\bar{\alpha}_{n}$ and $\bar{\beta}_{n}$ the geodesic sub-segments of $\bar{\gamma}_{n}$ meeting at $p$. Then replace $\bar{\beta}_{n}$ by its image under an isometry of $\mathbb{L}^{2}(\kappa)$ whose derivative maps the initial velocity vector of $\bar{\beta}_{n}$ to the terminal velocity vector of $\bar{\alpha}_{n}$, thereby straightening the concatenation of $\bar{\alpha}_{n}$ and $\bar{\beta}_{n}$ into an unbroken geodesic segment. We repeat this procedure at each breakpoint of $\bar{\gamma}_{n}$, and analogously for $\bar{\sigma}_{n}$. Concavity of the stacking implies that these unbroken geodesics will remain within the region bounded by the original stacking and therefore also have to meet. Moreover, since we applied an isometry in each step, the resulting unbroken geodesic segments retain the Lorentzian length of $\bar{\gamma}_{n}$ and $\bar{\sigma}_{n}$, respectively. Consequently, we have found two different geodesic segments in $\mathbb{L}^{2}(\kappa)$ that meet at $\bar{\tau}$-length strictly less than $D_{\kappa}$, a contradiction. ∎
@@ 4.2 Nonexistence of conjugate points under nonpositive timelike curvature
Cover $\gamma([0,1])$ by comparison neighborhoods $V_{1},\dots,V_{N}$ such that for a suitable partition $0=t_{0}<\dots<t_{N}=1$, $\gamma([t_{i},t_{i+1}])\subseteq V_{i+1}$ for all $i=0,\dots,N-1$. By (ii), there exists some $k_{0}$ such that also $\gamma_{p_{k}q_{k}}([t_{i},t_{i+1}])\subseteq V_{i+1}$ for all $k\geq k_{0}$ and all $i$. The proof of [7, Prop. 2.3] shows that then each $\gamma_{p_{k}q_{k}}|_{[t_{i},t_{i+1}]}$ as well as $\gamma|_{[t_{i},t_{i+1}]}$ are timelike maximizers. Since $\tau|_{V_{i}\times V_{i}}$ is continuous, we obtain
@@ Appendix A The Fréchet distance between curves
Let $\alpha\colon[0,1]\to[0,1]^{2}$ be a continuous, coordinate-wise non-decreasing map given by $\alpha(t)=(\varphi(t),\psi(t))$ with $\alpha(0)=(0,0)$ and $\alpha(1)=(1,1)$. Then there exists a coordinate-wise non-decreasing and $1$-Lipschitz continuous curve $\tilde{\alpha}=(\tilde{\varphi},\tilde{\psi})\colon[0,2]\to[0,1]^{2}$ that has the same image as $\alpha$.
@@
Finally, to show that $\tilde{\alpha}$ is $1$-Lipschitz, let $0\leq u<v\leq 2$. Then using monotonicity of $\theta$, $\varphi$ and $\psi$, hence of $\tilde{\varphi}$ and $\tilde{\psi}$,
@@
Let $N$ be the set of continuous, coordinate-wise non-decreasing maps $\alpha=(\varphi,\psi)\colon[0,1]\to[0,1]^{2}$ with $\alpha(0)=(0,0)$ and $\alpha(1)=(1,1)$. Then $\mathcal{N}\coloneqq\{\alpha([0,1])\mid\alpha\in N\}$ is compact with respect to the Hausdorff distance $d_{H}$.
@@
Denote by $L$ the set of all coordinate-wise non-decreasing $1$-Lipschitz maps $\eta\colon[0,2]\to[0,1]^{2}$ with $\eta(0)=(0,0)$ and $\eta(1)=(1,1)$. By the Theorem of Arzelà–Ascoli, $L$ is relatively compact in $C([0,2],[0,1]^{2})$. Since it is clearly also closed with respect to uniform convergence, it is indeed compact in $C([0,2],[0,1]^{2})$.
@@
Since $d_{H}(\iota(\eta),\iota(\zeta))\leq\sup_{u\in[0,2]}\|\eta(u)-\zeta(u)\|_{2}$, it follows that $\iota$ is continuous (indeed, $1$-Lipschitz). Thus $\mathcal{L}=\iota(L)$ is compact with respect to $d_{H}$. Finally, by Lemma A.1, $\mathcal{N}=\mathcal{L}$, concluding the proof. ∎
@@ Acknowledgments
This research was funded in part by the Austrian Science Fund (FWF) [Grant DOI 10.55776/EFP6 and 10.55776/J4913]. For open access purposes, the authors have applied a CC BY public copyright license to any author accepted manuscript version arising from this submission.